OpenStudy (anonymous):

A tire company manufactures 50 tires an hour with its current staff, the company can make 10 extra tires per hour for every extra worker it hires which equation represents this relationship mathematically, with x being the number of extra workers hired and y being the number of tires produced

OpenStudy (anonymous):

anyone?

OpenStudy (anonymous):

okay. assume it has C number of current staff. so C produce 50 tires in one hour.

OpenStudy (anonymous):

if y is the number of tires produced by each person on the staff, then Cy = 50 per hour.

OpenStudy (anonymous):

if the staff increases by 1, they produce 60 tires per houor, i.e if the number of people on staff is C+1, they produce 50 +10 tires per hour. so (C+1)y = 50 +1*10

OpenStudy (anonymous):

can you see where this is going?

OpenStudy (anonymous):

go on

OpenStudy (anonymous):

if they hire 2 extra people, (C+2)y = 50 +20 or (C+2)y = 50+2*10

OpenStudy (anonymous):

surely you can see what happens when they hire x new people?

OpenStudy (anonymous):

wait so c is x?

OpenStudy (anonymous):

no, C is the current number of people on the staff.

OpenStudy (anonymous):

if there are 5 people currently on the staff, and you hire two new people, what will your staff size be?

OpenStudy (anonymous):

ok there are choices

OpenStudy (anonymous):

its a mc question

OpenStudy (anonymous):

A. x=10+50y

OpenStudy (anonymous):

I gave you the logic behind the question. I am sure you can choose which answer is the correct one.

OpenStudy (anonymous):

Y=50+10X?

OpenStudy (anonymous):

yes. if Y is the total number of tires produced per hour.

OpenStudy (anonymous):

OK